# Moving Object Detection from Moving Camera Using Focus of Expansion Likelihood and Segmentation

### v1

@@ II RELATED WORKS
This section reviews prior efforts in moving object detection. There are numerous methods for detecting moving objects using static cameras. For instance, Rozumnyi et al. [4] detect fast-moving objects in their research. However, there is limited research on moving object detection from a moving camera.
Notable review papers on moving object detection include [5], and [6]. Based on these reviews, we have identified several key areas of research in moving object detection, including flow orientation-based methods, focus of expansion (FoE) based approaches, and adversarial network methods.
@@ III-A System Overview
Figure 2: System overview of the proposed method, FoELS.
@@
III-B Key Ideas to Overcome Challenges
In Section I, we identified four challenges for moving object detection from a moving camera. To address these challenges, FoELS integrates the following key ideas.
We incorporate an FoE-based approach, as it can address both straight-ahead camera motion and the misinterpretation of large optical flow from nearby static objects (Challenge 1). However, unlike [10], we do not assume a fixed FoE and instead allow it to vary with each frame. This allows us to handle various camera movements. To address the challenges of misinterpreting large optical flow from nearby static objects (Challenge 1) and insufficient flow in low-textured regions (Challenge 2), which [3] struggled with, we incorporate segmentation as an additional cue for detecting moving objects. Additionally, we introduce a method to address the challenge of ambiguity in parallel motion (Challenge 3), which will be explained in III-H. Finally, we add object-level refinement, which extracts moving objects after identifying moving pixels to address the challenge of detecting partially stationary objects (Challenge 4). The method will be explained in III-I. This paper’s primary objective is the detection of complete moving objects, rather than individual moving pixels, t o support downstream applications such as 3D reconstruction.
Ultimately, FoELS integrates optical flow and segmentation to overcome the limitations of previous methods.
@@ III-C System Details
First, frames $t-1$ and $t$ are used to compute the optical flow. Simultaneously, segmentation is performed on frame $t$, and each pixel is assigned a prior moving probability according to a manually predefined class-moving probability table. Sky regions identified through segmentation are removed, since optical flow cannot be computed there. Concurrently, static areas (e.g., ground, mountains, and buildings) are identified, and the flow within these regions is analyzed. If the flow existing ratio in the static area exceeds a specified threshold, the camera is considered to be in motion, and the FoE is computed using RANSAC. The inliers from the RANSAC process are attributed to camera motion, while the outliers are considered as moving pixels. Once the moving pixels are identified, we map them back to moving objects using the panoptic segmentation results.
@@
Figure 3: Detailed flowchart of the proposed method, with color coding corresponding to Fig. 2. The side image illustrates the process outlined in the flowchart.
@@ III-D Optical Flow Estimation
Accurate optical flow estimation forms the foundation of FoELS. We adopt the UniMatch model [12], which is trained on challenging benchmarks such as Sintel and KITTI, and achieves state-of-the-art accuracy. UniMatch provides dense flow maps, capturing subtle motion patterns critical for subsequent analysis.
@@ III-E Segmentation
Segmentation is a widely applicable and extensively investigated area in computer vision. Numerous studies have focused on segmentation [13, 14, 15, 16, 17, 18, 19]. Common semantic segmentation approaches proved insufficient for detecting moving objects, as it is crucial to distinguish between individual object instances, some of which may be moving while others remain static. Therefore, panoptic segmentation, which combines semantic and instance segmentation, was adopted for this work.
Several segmentation models capable of handling panoptic segmentation tasks have been proposed recently, such as Mask2Former [20]. Among these advanced models, OneFormer [21] is currently the state-of-the-art panoptic segmentation model. Consequently, the OneFormer panoptic segmentation model is utilized in this approach. Each class is assigned a prior probability reflecting its tendency to be dynamic. For instance, sky and building classes have low moving probabilities, while vehicle and pedestrian classes have higher values.
@@ III-H Moving Pixel Probability Estimation
$d_{a}$ $=\arccos\left(\frac{\mathbf{v}_{F}\cdot\mathbf{v}_{P}}{||\mathbf{v}_{F}||\cdot||\mathbf{v}_{P}||}\right),$ (5)
$d_{l}$ $=||\mathbf{v}_{P}||/\overline{||\mathbf{v}_{P,static}||},$ (6)
where, $\mathbf{v}_{F}$ is the vector from the FoE to the point, and $\mathbf{v}_{P}$ is the optical flow vector at the point. The term $d_{a}$ represents the angular difference calculated from these vectors. $d_{l}$ is the relative flow length difference, by $\overline{||\mathbf{v}_{P,static}||}$, which denotes the mean optical flow magnitude observed in static regions.
@@ III-I Object-Level Refinement
Finally, the computed moving pixel probabilities are aggregated to an object level. This step is crucial for ensuring that an entire object is classified as moving, even if only a portion of it exhibits detectable motion (addressing Challenge 4). To achieve this, a binary moving pixel mask is first generated by thresholding the posterior moving pixel probability $P_{M}^{\prime}$. A threshold of $0.250.5^{2}=$ 0.25 is used for $P_{M}^{\prime}$, reflecting the fact that $P_{M}^{\prime}$ is a product of two probabilities ($P_{seg}$ and $P_{FoE}$); this threshold implies that both contributing probabilities are at least 0.5. Subsequently, an object-level moving mask is derived. For each object instance identified by the panoptic segmentation, the percentage of pixels within that instance that are marked as moving in the binary pixel mask is calculated. If this percentage exceeds a threshold of 0.01, the entire object instance is classified as moving. This low threshold is employed to effectively detect objects where only a small part is in motion, such as the tail of an animal or a limb of a person, which can sometimes constitute as little as approximately 3% of the total object area.
@@ IV-A2 Qualitative Evaluation Dataset
To assess the performance of FoELS in real-world conditions, a custom traffic video dataset was captured. This dataset encompasses a variety of challenging scenarios, including vehicles moving parallel to the camera, stopped vehicles, and operation in low-textured environments. Furthermore, it uniquely features instances of camera zooming, a condition not typically represented in standard moving object detection datasets such as DAVIS 2016 and FBMS-59.
@@ IV-C Results
Fig. 4 shows an example of the visual results from the above evaluation. This figure illustrates the step-by-step results of the process detailed in Section III. The first row displays: (a) the input frame, (b) the segmentation result, where different colors denote distinct classes, and (c) the prior moving probability derived from segmentation. The prior probability is visualized using a jet colormap, where red indicates higher probability and blue signifies lower probability. The second row presents: (d) the optical flow, with orientation encoded by color, (e) the optical flow field highlighting Focus of Expansion (FoE) inliers (green arrows) and outliers (red arrows). An existing FoE in the image is marked with a thick red cross. (f) The FoE-based moving likelihood, also depicted using a jet colormap. The third row shows: (g) the posterior moving pixel probability, calculated as the product of the prior moving probability and the FoE-based moving likelihood, (h) the refined object-level moving mask, demonstrating the aggregation of moving pixels to an object level, and (i) the final moving object mask overlaid on the input image. In this particular example, the bear’s hand remains stationary while the bear is walking. Nevertheless, FoELS successfully extracts the entire bear due to the object-level refinement process.
@@
Fig. 5 compares the results of the Adversarial Network with those of FoELS for the same scene as Fig. 1. The left side shows the results of the Adversarial Network, while the right side displays the results of FoELS. The Adversarial Network falsely detects nearby vegetation and poles as moving objects due to their significantly different optical flow compared to the background. In contrast, FoELS successfully identifies only the genuinely moving objects by primarily relying on FoE-based flow orientation analysis.
Figure 5: Comparison results with Adversarialnet (left) and FoELS (right).
Fig. 6 shows the step by step visualization results of the same scene as Fig. 1. In this example, it can be seen why FoELS can correctly detect cars almost moving parallel to the camera, and the nearby static pole, despite exhibiting large optical flow, is correctly identified as stationary.
Figure 6: Example visual result of FoELS on custom traffic video. Each image corresponds to the same visualization format as shown in Fig. 4.
Fig. 7 compares the results of the Adversarial Network with those of FoELS on the custom zoom in/out video evaluation, where the camera remains stationary while zooming. The left panel illustrates the results of the Adversarial Network, while the right panel displays the results of FoELS. In this initial frame, the camera is nearly stationary, and no zoom is applied. However, the Adversarial Network produces numerous false positives. This indicates a lack of robustness in the Adversarial Network when applied to novel, untrained scenes. Conversely, FoELS exhibits no false positives in this scenario.
Figure 7: Comparison results: Adversarial Network (left) and FoELS (right) on a custom zoom in/out video (initial frame).
Fig. 8 presents a similar comparison between the Adversarial Network and FoELS. In this instance, the camera is actively zooming in while the train is in motion. An incoming train is positioned near the center of the image, while simultaneously the background exhibits motion due to the camera zoom. Notably, the Adversarial Network fails to detect any moving objects. In contrast, FoELS successfully identifies the approaching train while correctly disregarding the background motion induced by the zoom.
Figure 8: Comparison results: Adversarial Network (left) and FoELS (right) on a custom zoom in/out video (train approaching during zoom).
Fig. 9 illustrates the intermediate processing steps for the same frame of Fig. 8, employing the visualization format detailed in Fig. 4. This visualization demonstrates the successful detection of the moving train by FoELS.
Figure 9: Zoom in/out scene visual result of FoELS on the same frame with Fig. 8. Each image corresponds to the same visualization format as shown in Fig. 4.
@@
Fig. 10 (b) illustrates an example of the camel scene in the DAVIS 2016 dataset. In this scene, it successfully detects only the moving camel while ignoring the static camel.
Other scenes also contain challenging scenarios, where some potential moving objects are in motion while others remain static. However, FoELS successfully detects the moving objects in all of them.
@@
((b)) DAVIS:camel
((c)) DAVIS:dance-twiri
((d)) Custom:truck
((e)) FBMS:cars4
((f)) FBMS:tennis
Figure 10: Example visual results of FoELS. Each image corresponds to the same visualization format as shown in Fig. 4.
@@ V CONCLUSION
Future work will focus on optimizing FoELS for real-time performance, enhancing camera motion detection (e.g., via direct estimation using deep neural networks as an alternative to FoE-based inference), and integrating object tracking (e.g. [24, 25]) to leverage temporal information for improved accuracy and consistency.

### v2

@@ II RELATED WORKS
This section reviews prior efforts in optical flow estimation, segmentation, and moving object detection, which form the foundation of our approach.
II-A Optical Flow Estimation
The field of optical flow estimation has evolved significantly with deep learning approaches. CNN-based approaches [4, 5] demonstrated end-to-end optical flow prediction. RAFT [6] introduced 4D correlation volumes and GRU [7] refinement, dramatically improving accuracy. Recent transformer-based approaches [8, 9, 10, 11] including UniMatch [12] and MemFlow [13] achieve state-of-the-art accuracy.
II-B Segmentation
Segmentation is extensively studied in computer vision [14, 15, 16, 17, 18, 19, 20]. Panoptic segmentation, which combines semantic and instance segmentation, is essential to distinguish between individual object instances. Among recent models [21], OneFormer [22] is currently the state-of-the-art panoptic segmentation model.
II-C Moving Object Detection
There are numerous methods for detecting moving objects using static cameras. For instance, Rozumnyi et al. [23] detect fast-moving objects in their research. However, there is limited research on moving object detection from a moving camera.
Notable review papers on moving object detection include [24], and [25]. Based on these reviews, we have identified several key areas of research in moving object detection, including flow orientation-based methods, focus of expansion (FoE) based approaches, and adversarial network methods. The FoE is the point from which optical flow vectors radiate when the camera moves forward [26]. Moving objects exhibit flow patterns deviating from this radial structure, making FoE-based analysis effective for motion detection.
@@ III-A System Overview
III-B Contributions and Key Ideas to Overcome Challenges
In Section I, we identified four challenges for moving object detection from a moving camera.
To address these challenges, FoELS integrates the following key ideas, representing three distinct contributions:
Contribution 1: Introduction of macroscopic perspective. Previous approaches all focus on microscopic information (optical flow). We incorporate an FoE-based approach for microscopic pixel-level analysis, which addresses both straight-ahead camera motion and the misinterpretation of large optical flow from nearby static objects (Challenge 1). However, unlike [30], we do not assume a fixed FoE and instead allow it to vary with each frame, enabling us to handle various camera movements. Critically, we first introduce macroscopic information (texture information through segmentation) to the moving object detection problem, addressing the challenges of misinterpreting large optical flow from nearby static objects (Challenge 1) and insufficient flow in low-textured regions (Challenge 2), which [3] struggled with. Additionally, we introduce object-level refinement to extract complete moving objects even when only parts exhibit motion (Challenge 4), explained in III-I. This leverages panoptic segmentation to ensure detection of complete objects rather than individual moving pixels.
Contribution 2: Probabilistic integration framework. We do not integrate naively by taking AND or OR operations. Instead, we probabilistically integrate microscopic (optical flow) and macroscopic (segmentation) perspectives. This probabilistic combination enables our algorithm to achieve high accuracy robustly.
Contribution 3: Original improvement for parallel motion detection. Through experimentation, we discovered that probabilistic integration alone is insufficient for parallel motion scenarios, which previous methods also cannot solve. We therefore developed a novel solution: incorporating optical flow length consideration into the FoE-based likelihood (Challenge 3), detailed in III-H.
@@ III-C System Details
First, frames t1t-1 and $t$ are used to compute the optical flow. Simultaneously, segmentation is performed on frame $t$, and each pixel is assigned a prior moving probability according to a manually predefined class-moving probability table. Sky regions identified through segmentation are removed, since optical flow cannot be computed there. Concurrently, static areas (e.g., ground, mountains, and buildings) are identified, and the flow within these regions is analyzed. If the flow existing ratio in the static area exceeds a specified threshold, the camera is considered to be in motion, and the FoE is computed using RANSAC. The inliers from the RANSAC process are attributed to camera motion, while the outliers are considered as moving pixels. Once the moving pixels are identified, we map them back to moving objects using the panoptic segmentation results.
@@
Figure 2: Detailed flowchart of the proposed method. The side image illustrates the process outlined in the flowchart.
@@ III-D Optical Flow Estimation
Based on our research in Section II and Sintel benchmark rankings, we identified UniMatch [12] and MemFlow [13] as top candidates with publicly available code. Quantitative comparisons using the DAVIS 2016 dataset (TABLE IV) led us to select UniMatch, which provides dense flow maps capturing subtle motion patterns critical for subsequent FoE-based analysis.
@@ III-E Segmentation
As discussed in Section II, we utilize the OneFormer panoptic segmentation model, which is currently the state-of-the-art approach. In FoELS, each class is assigned a prior probability reflecting its tendency to be dynamic. For instance, sky and building classes have low moving probabilities, while vehicle and pedestrian classes have higher values.
@@ III-H Moving Pixel Probability Estimation
[…]
@@ III-I Object-Level Refinement
Finally, the computed moving pixel probabilities are aggregated to an object level. This step is crucial for ensuring that an entire object is classified as moving, even if only a portion of it exhibits detectable motion (addressing Challenge 4). To achieve this, a binary moving pixel mask is first generated by thresholding the posterior moving pixel probability $P_{M}^{\prime}$. A threshold of $0.5^{2}=0.25$ is used for $P_{M}^{\prime}$, reflecting the fact that $P_{M}^{\prime}$ is a product of two probabilities ($P_{seg}$ and $P_{FoE}$); this threshold implies that both contributing probabilities are at least 0.5. Subsequently, an object-level moving mask is derived. For each object instance identified by the panoptic segmentation, the percentage of pixels within that instance that are marked as moving in the binary pixel mask is calculated. If this percentage exceeds a threshold of 0.01, the entire object instance is classified as moving. This low threshold is employed to effectively detect objects where only a small part is in motion, such as the tail of an animal or a limb of a person, which can sometimes constitute as little as approximately 3% of the total object area.
@@ IV-A2 Qualitative Evaluation Dataset
To assess the performance of FoELS in real-world conditions, a custom traffic video dataset was captured. This dataset was specifically designed to address failure modes of existing methods and evaluate FoELS’s contributions. The chosen scenarios are: (1) Parallel-moving vehicles: The most challenging scenario where flow direction of moving objects closely aligns with background, requiring our novel flow length consideration approach. (2) Stationary vehicles: Common in traffic scenarios, testing the ability to distinguish truly moving objects from temporarily stopped ones. (3) Low-textured environments: Challenging for optical flow-based methods, demonstrating the value of integrating segmentation as a macroscopic cue. (4) Camera zoom: Not represented in standard datasets (DAVIS 2016, FBMS-59), testing the robustness of FoE-based approach to non-translational camera motion. Additionally, we added experiments on opposite-direction, cross-direction motion, and crowded scenes to demonstrate FoELS’s general applicability beyond these specific scenarios.
@@ IV-C Results
Fig. 3 shows an example of the visual results from the above evaluation. This figure illustrates the step-by-step results of the process detailed in Section III. The first row displays: (a) the input frame, (b) the segmentation result, where different colors denote distinct classes, and (c) the prior moving probability derived from segmentation. The prior probability is visualized using a jet colormap, where red indicates higher probability and blue signifies lower probability. The second row presents: (d) the optical flow, with orientation encoded by color, (e) the optical flow field highlighting FoE inliers (green arrows) and outliers (red arrows). An existing FoE in the image is marked with a thick red cross. (f) The FoE-based moving likelihood, also depicted using a jet colormap. The third row shows: (g) the posterior moving pixel probability, calculated as the product of the prior moving probability and the FoE-based moving likelihood, (h) the refined object-level moving mask, demonstrating the aggregation of moving pixels to an object level, and (i) the final moving object mask overlaid on the input image. In this particular example, the bear’s hand remains stationary while the bear is walking. Nevertheless, FoELS successfully extracts the entire bear due to the object-level refinement process.
@@
Fig. 4 (a) compares the results of the Adversarial Network with those of FoELS for the same scene as Fig. 1. The left side shows the results of the Adversarial Network, while the right side displays the results of FoELS. The Adversarial Network falsely detects nearby vegetation and poles as moving objects due to their significantly different optical flow compared to the background. In contrast, FoELS successfully identifies only the genuinely moving objects by primarily relying on FoE-based flow orientation analysis.
((a)) Traffic
((b)) Zoom:initial
((c)) Zoom:train
Figure 4: Comparison results with AdversarialNet (left) and FoELS (right) across different scenarios. AdversarialNet exhibits limited generalization to unseen scenes, while FoELS maintains robust performance without scene-specific tuning. The dramatic visual improvement reflects the difference between real-world complexity and standard datasets.
Fig. 5 (b) shows the step by step visualization results of the same scene as Fig. 1. In this example, it can be seen why FoELS can correctly detect cars almost moving parallel to the camera, and the nearby static pole, despite exhibiting large optical flow, is correctly identified as stationary.
Fig. 4 (b) compares the results of the Adversarial Network with those of FoELS on the custom zoom in/out video evaluation, where the camera remains stationary while zooming. The left panel illustrates the results of the Adversarial Network, while the right panel displays the results of FoELS. In this initial frame, the camera is nearly stationary, and no zoom is applied. However, the Adversarial Network produces numerous false positives. This indicates a lack of robustness in the Adversarial Network when applied to novel, untrained scenes. Conversely, FoELS exhibits no false positives in this scenario.
Fig. 4 (c) presents a similar comparison between the Adversarial Network and FoELS. In this instance, the camera is actively zooming in while the train is in motion. An incoming train is positioned near the center of the image, while simultaneously the background exhibits motion due to the camera zoom. Notably, the Adversarial Network fails to detect any moving objects. In contrast, FoELS successfully identifies the approaching train while correctly disregarding the background motion induced by the zoom.
Fig. 5 (c) illustrates the intermediate processing steps for the same frame of Fig. 4 (c), employing the visualization format detailed in Fig. 3. This visualization demonstrates the successful detection of the moving train by FoELS.
@@
These scenes contain challenging scenarios, where some potential moving objects are in motion while others remain static. However, FoELS successfully detects the moving objects in all of them.
@@
((b)) Custom:traffic
((c)) Custom:zoom
((d)) Custom:parallel
((e)) Custom:cross
((f)) Custom:opposite in crowded
Figure 5: Example visual results of FoELS on various motion types including parallel, opposite-direction, cross-direction, and crowded scenes. See Fig. 3 for the 9-subimage format.
@@ IV-D Ablation Studies
We also compared different optical flow methods to select the most suitable one for FoELS. TABLE IV shows the quantitative comparison between MemFlow and UniMatch.
@@
TABLE IV: Comparison of optical flow methods. The values represent the average IoU scores of FoELS over the DAVIS 2016 train-val-movobj sequences using different optical flow methods.
Optical Flow Method IoU
MemFlow 0.736
UniMatch (FoELS) 0.757
@@ V CONCLUSION
V-A Limitations and Future Work
FoELS achieves strong accuracy but is computationally intensive due to the trade-off between accuracy and real-time feasibility. Temporal instability also occurs from frame-by-frame processing. Future work includes optimization strategies: model pruning, quantization, lighter architectures, parallel processing, and adding tracking.